OpenStudy (anonymous):

Which explains why the sequence 64, 4, 1.4, ... is arithmetic or geometric? Is it geometric or artihmetic? decreases by which factor? 1/16 or 4? I think geometric and decreaes by 4. Help please? @bibby

OpenStudy (anonymous):

@Tazmaniadevil

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

The y intercept is 8 so mark that point on the y axis. The slope is negative 1 over 4. That means the rise is -1 and the run is positive 4. From the y intercept of 8, follow the direction given by the slope and go down one unit and right 4 units. Mark that point on the graph. Since two points determine a line, draw the line through the two points and put arrows on the ends because a line goes on and on forever in both directions

OpenStudy (anonymous):

oh thank you!! Ok, I will read what you wrote. Just sec

OpenStudy (anonymous):

Ok so I think I got it right? The sequence is geometric because it decreases by a factor of 4. ?

OpenStudy (anonymous):

@josgilbertsen

OpenStudy (campbell_st):

well the sequence given isn't geometric or arithmetic... can you please check its not 64, 4, 1/4, ..... if it is, then its geometric with a common ratio of 1/16... as you said...

OpenStudy (anonymous):

It is asking if it is geometric or arithmetic in sequence.

OpenStudy (anonymous):

dang i got it wrong. Ok i think I understand it.

OpenStudy (campbell_st):

look at it this way 64 x 1/16 = 4 4 x 1/16 = 1/4 in geometric sequences... you are multiplying... the previous term...

OpenStudy (anonymous):

ok I am writing that down in my notes. Thank you. Can I ask you another question? It is about variables.

OpenStudy (campbell_st):

well if you post it in the normal way I'll look at it

OpenStudy (anonymous):

yes I will definetly!

OpenStudy (anonymous):

geometric because it decreases by a factor of 1/16